# Non-perturbative self-consistent electron-phonon spectral functions and transport

### v1

@@ Acknowledgements.
We thank Neil R. Wilson and Vladimir I. Fal’ko for kindly providing the data for Ref. [73]. S.P. is a Research Associate of the Fonds de la Recherche Scientifique - FNRS. This work was supported by the Fonds de la Recherche Scientifique - FNRS under Grants number T.0183.23 (PDR) and T.W011.23 (PDR-WEAVE). This publication was supported by the Walloon Region in the strategic axe FRFS-WEL-T. Computational resources have been provided by the EuroHPC JU award granting access to MareNostrum5 at Barcelona Supercomputing Center (BSC), Spain (Project ID: EHPC-EXT-2023E02-050), and by the Consortium des Équipements de Calcul Intensif (CÉCI), funded by the FRS-FNRS under Grant No. 2.5020.11 and computational resources on Lucia, the Tier-1 supercomputer of the Walloon Region with infrastructure funded by the Walloon Region under the grant agreement n°1910247.References
@@
- Poncé et al. [2020] S. Poncé, W. Li, S. Reichardt, and F. Giustino, First-principles calculations of charge carrier mobility and conductivity in bulk semiconductors and two-dimensional materials, Rep. Prog. Phys. 83, 036501 (2020).
@@
- Note [1] See Supplemental Material [URL will be inserted by publisher], which includes Refs. [98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114], for the computational details and additional results.
@@
- Lihm and Poncé [2024] J.-M. Lihm and S. Poncé, Non-perturbative self-consistent electron-phonon spectral functions and transport [link to be added], Materials Cloud Archive (2024).

### v2

@@ Acknowledgements.
We thank Neil R. Wilson and Vladimir I. Fal’ko for kindly providing the data for Ref. [73], and the anonymous referees who provided invaluable insights which led to an improved version of this work. S.P. is a Research Associate of the Fonds de la Recherche Scientifique - FNRS. This work was supported by the Fonds de la Recherche Scientifique - FNRS under Grants number T.0183.23 (PDR) and T.W011.23 (PDR-WEAVE). This publication was supported by the Walloon Region in the strategic axe FRFS-WEL-T. Computational resources have been provided by the EuroHPC JU award granting access to MareNostrum5 at Barcelona Supercomputing Center (BSC), Spain (Project ID: EHPC-EXT-2023E02-050), and by the Consortium des Équipements de Calcul Intensif (CÉCI), funded by the FRS-FNRS under Grant No. 2.5020.11 and computational resources on Lucia, the Tier-1 supercomputer of the Walloon Region with infrastructure funded by the Walloon Region under the grant agreement n°1910247.References
@@
- Shishkin and Kresse [2007] M. Shishkin and G. Kresse, Self-consistent calculations for semiconductors and insulators, Phys. Rev. B 75, 235102 (2007).
- Note [1] See Supplemental Material [URL will be inserted by publisher], which includes Refs. [107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125], for the derivation of the formula, computational details, and additional results.
@@
- Hybertsen and Louie [1986] M. S. Hybertsen and S. G. Louie, Electron correlation in semiconductors and insulators: Band gaps and quasiparticle energies, Phys. Rev. B 34, 5390 (1986).
@@
- Poncé et al. [2014] S. Poncé, G. Antonius, Y. Gillet, P. Boulanger, J. Laflamme Janssen, A. Marini, M. Côté, and X. Gonze, Temperature dependence of electronic eigenenergies in the adiabatic harmonic approximation, Phys. Rev. B 90, 214304 (2014).
@@
- Kas et al. [2014] J. J. Kas, J. J. Rehr, and L. Reining, Cumulant expansion of the retarded one-electron Green function, Phys. Rev. B 90, 085112 (2014).
- Kas and Rehr [2017] J. J. Kas and J. J. Rehr, Finite temperature Green’s function approach for excited state and thermodynamic properties of cool to warm dense matter, Phys. Rev. Lett. 119, 176403 (2017).
@@
- Capone and Ciuchi [2003] M. Capone and S. Ciuchi, Polaron crossover and bipolaronic metal-insulator transition in the half-filled Holstein model, Phys. Rev. Lett. 91, 186405 (2003).
- Bauer et al. [2011] J. Bauer, J. E. Han, and O. Gunnarsson, Quantitative reliability study of the Migdal-Eliashberg theory for strong electron-phonon coupling in superconductors, Phys. Rev. B 84, 184531 (2011).
- Franchini et al. [2021] C. Franchini, M. Reticcioli, M. Setvin, and U. Diebold, Polarons in materials, Nat. Rev. Mater. 10.1038/s41578-021-00289-w (2021).
- Sze and Ng [2006] S. M. Sze and K. K. Ng, Physics of Semiconductor Devices (John Wiley & Sons, Inc., Hoboken, NJ, USA, 2006).
- Lundstrom [2010] M. Lundstrom, Fundamentals of Carrier Transport, 2nd ed. (Cambridge University Press, Cambridge, 2010).
@@
- Lihm and Poncé [2025] J.-M. Lihm and S. Poncé, Non-perturbative self-consistent electron-phonon spectral functions and transport, Materials Cloud Archive 10.24435/materialscloud:k1-rt (2025).
- Dunn [1975] D. Dunn, Electron-phonon interactions in an insulator, Can. J. Phys. 53, 321 (1975).
- Nery and Mauri [2022] J. P. Nery and F. Mauri, Non-perturbative Green’s function method to determine the electronic spectral function due to electron-phonon interactions: Application to a graphene model from weak to strong coupling, ArXiv220311289 Cond-Mat (2022).
@@
End Matter
Comparison of the self-energies— Figure 4 compares the and sc self-energies at 100 K and 300 K. The one-shot self-energy shows a sharp peak in the imaginary part at , where is the LO phonon frequency. This peak, associated with phonon emission, is significantly damped and smeared in the sc self-energy due to the self-consistent broadening of the spectral function. This self-consistent regularization becomes more pronounced at higher temperatures.
At 300 K, the sc self-energy reaches around 0.1 eV. While this value is smaller than the full bandwidth of the valence band of InSe (1.2 eV, see Fig. LABEL:supp-fig:InSe_itp of the SM [72]), it is larger than the width of the Mexican-hat feature at the valence band edge (0.07 eV). Thus, the self-energy leads to a significant renormalization near the valence band edge, qualitatively altering the spectral function.
Tail of the spectral functions— Compared to the and cumulant spectral functions, the sc spectral functions show a clear qualitative difference in insulators and nondegenerate semiconductors: the former exhibits an unphysically long tail in the mid-gap region, which is absent in the latter. Figures 5(a, b) show the density of states (DOS)
and the hole DOS (). In Fig. 5(a), the DOS from the and the cumulant methods show a slow decay (gray shaded curves). This is a result of the Lorentzian broadening of the bare Green’s function (see Sec. LABEL:supp-sec:decay of the SM [72] for a detailed derivation.) Importantly, reducing the artificial broadening reduces the amplitude of the tail, but does not change the scaling. When multiplied by the hole occupation factor, which increases exponentially with energy up to the chemical potential, the hole DOS continues to increase above the renormalized band edge, peaking at the chemical potential [Fig. 5(b)]. As a result, in nondegenerate semiconductors, the and cumulant methods incorrectly predict an unphysically large amount of carriers in the mid-gap region, where the spectral function is small. In contrast, the sc spectral function decays faster than (red shaded curve), and the hole DOS peaks at the renormalized band edge as expected.
Doping dependence of the conductivity— Figure 5(c) shows the doping dependence of the hole mobility of monolayer InSe at . Surprisingly, the (green plus markers) and cumulant mobility (blue squares) show significant changes as a function of hole concentration. This result contrasts sharply with what is typically observed in nondegenerate semiconductors with dominant -ph scattering [92, 93]. We attribute this artifact to the unphysical tail in the spectral functions, which places most of the holes near the chemical potential, where the actual spectral function is small. Due to this strong doping dependence, the intrinsic (low doping limit) mobility cannot be properly calculated using the and cumulant approximations. To alleviate this problem, we developed a truncation scheme (see Eq. (LABEL:supp-eq:trunc_occ_p_doped) of the SM [72]) that removes the tail of the spectral function beyond the local minima in the hole DOS. Then, we obtain a doping-independent Bubble- mobility (green crosses). However, this truncation could not be applied to the cumulant spectral functions of InSe due to its large broadening.
In contrast, the sc mobility (orange circles) as well as the BTE mobility (blue and red triangles) remain doping-independent in the nondegenerate regime (white background). Doping dependence is observed only in the metallic regime (gray background), where the chemical potential lies within the valence band. Therefore, it is highly desirable to use the sc method, rather than the or cumulant approximations, for studying transport in nondegenerate semiconductors.